# Iron Sharpens Iron: Defending Against Attacks in Machine-Generated Text Detection with Adversarial Training

### v1

@@ A.1 Dataset
We used the Semeval Task8 dataset Siino (2024) as the primary dataset for training the detector. This large-scale dataset includes MGTs from ChatGPT 4, Davinci, Bloomz, Dolly, and Cohere, as well as HWTs from WikiHow, Reddit, arXiv, Wikipedia, and PeerRead. Moreover, the average token length of the dataset is 623.2521. For each scenario, we employed distinct datasets for testing, as detailed in Tabel 3.
@@ A.3 Evaluation Metrics
For detector, a lower ASR $\downarrow$ indicates better defense performance, while for adversary, a higher ASR $\uparrow$ signifies a stronger attack effectiveness.
@@
Perturbation Rate ($\mathrm{Pert.}$). $\mathrm{Pert.}$ measures the lexical difference between the adversarial text and the original text. It is defined as the ratio of the number of perturbed tokens to the total number of tokens in the text. A lower perturbation rate $\downarrow$ indicates that the adversarial text remains more similar to the original text.
@@
where $N$ is the total number of tokens, $w_{i}$ represents the $i$-th token, and $P(w_{i}|w_{1},w_{2},\dots,w_{i-1})$ is the probability assigned by the language model. Generally, a lower PPL variation $\downarrow$ indicates that the quality of the adversarial text is closer to that of the original text. We use GPT-2 Radford et al. (2019) to compute PPL.
@@
where $\mathbf{E}_{\text{orig}}$ and $\mathbf{E}_{\text{adv}}$ are the sentence embeddings of the original and adversarial texts, respectively, generated by the USE model, and $\|\mathbf{E}\|$ denotes the Euclidean norm. A higher USE score $\uparrow$ indicates greater semantic similarity. We use USE (Universal Sentence Encoder) to calculate USE score.
@@
where $N_{\text{words}}$ is the total number of words in the text, $N_{\text{sentences}}$ is the total number of sentences, and $N_{\text{syllables}}$ is the total number of syllables. A smaller $\Delta r$downarrow indicates less change in readability.
@@ A.4.1 Defense Methods
Edit Pretraining (EP) Wang et al. (2024b): A data augmentation method that blends Mixed Edit Attack into the training set.
Paraphrasing Pretraining (PP) Wang et al. (2024b): A data augmentation method that blends Paraphrasing Attack into the training set.
@@
CERT-ED Huang et al. (2024): An adversarial training method that adapts randomized deletion to effectively safeguard natural language classification models from diverse edit-based adversarial operations, including synonym substitution, insertion, and deletion.
RanMask Zeng et al. (2023): An adversarial training method that randomly masks a proportion of words in the input text, thereby mitigating both word- and character-level adversarial perturbations without assuming prior knowledge of the adversaries’ synonym generation.
Text-RS Zhang et al. (2024b): An adversarial training method that treats discrete word substitutions as continuous perturbations in the embedding space to reduce the complexity of searching through large discrete vocabularies and bolster the model’s robustness.
Text-CRS Zhang et al. (2024a): An adversarial training method that is built on randomized smoothing, encompassing various word-level adversarial manipulations—such as synonym substitution, insertion, deletion, and reordering—by modeling them in both embedding and permutation spaces.
@@ A.4.2 Attack Methods
HMGC Zhou et al. (2024): A black-box zero-query adversarial attack framework. It uses a surrogate model to approximate the detector, ranks words by gradient sensitivity and PPL, and replaces high-importance words via an encoder-based masked language model. Constraints ensure fluency and semantic consistency, while dynamic adversarial learning refines the attack strategy.
@@ A.5 Detailed Information of Text Perturbation Method
In this section, we introduce 9 Text Perturbation Methods mentioned in Table 1. Detailed information of Adversarial Attack Methods can be found in Appendix A.4.2.
@@
Appendix B Experience on Defense
@@ B.1 Defense Performance under Different Attack Strengths
B.2 Defense Adaptation to Different Backbones
Model Metric Baseline CERT-ED Greater-D
ALBERT Base (12M) ASR(%$\downarrow$ 63.58 46.71 35.62
ALBERT Large (18M) ASR(%$\downarrow$ 97.14 45.45 43.40
DeBERTa Base (86M) ASR(%$\downarrow$ 22.95 29.37 4.10
DeBERTa Large (304M) ASR(%$\downarrow$ 51.72 25.07 10.25
RoBERTa Base (125M) ASR(%$\downarrow$ 81.88 26.78 15.32
RoBERTa Large (355M) ASR(%$\downarrow$ 30.21 34.66 5.38
XLM-RoBERTa Large (561M) ASR(%$\downarrow$ 81.99 40.04 1.73
Appendix C Ablation Study
@@
Appendix D Mathematical Analysis of Perturbation Rate and Query Complexity
@@
D.1 Perturbation Rate Analysis
@@
D.2 Query Complexity Analysis
@@
Appendix E Adversarial Training Process
@@
Appendix F Case Study
@@
Method Text Result
Original MGT Suggested by the Scottish Parliamentary Constituencies Commission in 2003, and adopted at Holyrood on 1 May 2004. Area of Scotland’s 32nd largest council area - covering parts of East Renfrewshire Council Area Machine-written [Uncaptioned image]
PWWS Suggested by the Scottish Parliamentary Constituencies Commission in 2003, and adoptions at Holyrood on 1 Probability 2004. Area of Scotland’s 32nd highest council area - covering item of East Renfrewshire Council Area Human-written (Succeeded) [Uncaptioned image]
TextFooler Suggested by the Scottish Parliamentary Constituencies Commission in 2003, and adopted at Holyrood on 1 May 2004. Area of Scotland’s 32nd largest council area - covering parts of East Renfrewshire Council Area Machine-written (Failed) [Uncaptioned image]
BERTAttack Suggested by the Scottish rural Constituencies Commission in 2003, and abolished at Holyrood on 1 May 2004. Area of ward 32nd most council constituency - almost all of East Renfrewshire Council Area Human-written (Succeeded) [Uncaptioned image]
A2T Suggested by the Scottish Parliamentary Constituencies Commission in 2003, and adopted at Holyrood on 1 May 2004. Area of Scotland’s 32nd largest council area - covering parts of East Renfrewshire Council Area Machine-written (Failed) [Uncaptioned image]
FastTextDodger Suggested by the Scottish Parliamentary Constituencies Commission in 2003: and adopted at Holyrood on 1 May 2004. Area of’ ’s 32nd largest council area - covering parts of East council Council Human-written (Succeeded) [Uncaptioned image]
ABP Suggested aside the Scottish Parliamentary Constituencies Commission indium 2003, and adopted atomic number 85 Holyrood on ace Crataegus oxycantha 2004. Area of Scotland’s 32nd largest council area - covering parts of East Renfrewshire Council Area Human-written (Succeeded) [Uncaptioned image]
HQA Suggested by the Scottish Parliamentary Constituencies Commission in 2003, and adopted at Holyrood on 1 May 2004. Area of Scotland’s 2004. sphere council area thirty-second prominent council of East Renfrewshire Council Area Human-written (Succeeded) [Uncaptioned image]
T-PGD Introduced. by. Scottish Parliamentary Resituiances Commission in 2002, and adopted at Holyrood on 1 May 2004. All of Scotland’s 32 The largest councils area - covering parts of East Renfrewshire Council Area Human-written (Succeeded) [Uncaptioned image]
Greater-A Suggested by the Grave Parliamentary Constituencies Commission in 2003, and adopted at Holyrood on 1 May 2004. Area Scotland 32nd biggest council area - covering parts of East Renfrewshire Council Area Human-written (Succeeded) [Uncaptioned image]

### v2

@@ A.1 Dataset
We used the Semeval Task8 dataset Siino (2024) as the primary dataset for training the detector. This large-scale dataset includes MGTs from ChatGPT 4, Davinci, Bloomz, Dolly, and Cohere, as well as HWTs from WikiHow, Reddit, arXiv, Wikipedia, and PeerRead. Moreover, the average token length of the dataset is 623.2521. For each scenario, we employed distinct datasets for testing, as detailed in Tabel 5.
@@ A.3 Evaluation Metrics
For detector, a lower ASR indicates better defense performance, while for adversary, a higher ASR signifies a stronger attack effectiveness.
@@
Perturbation Rate ($\mathrm{Pert.}$). $\mathrm{Pert.}$ measures the lexical difference between the adversarial text and the original text. It is defined as the ratio of the number of perturbed tokens to the total number of tokens in the text. A lower perturbation rate indicates that the adversarial text remains more similar to the original text.
@@
where $N$ is the total number of tokens, $w_{i}$ represents the $i$-th token, and $P(w_{i}|w_{1},w_{2},\dots,w_{i-1})$ is the probability assigned by the language model. Generally, a lower PPL variation indicates that the quality of the adversarial text is closer to that of the original text. We use GPT-2 Radford et al. (2019) to compute PPL.
@@
where $\mathbf{E}_{\text{orig}}$ and $\mathbf{E}_{\text{adv}}$ are the sentence embeddings of the original and adversarial texts, respectively, generated by the USE model, and $\|\mathbf{E}\|$ denotes the Euclidean norm. A higher USE score indicates greater semantic similarity. We use USE (Universal Sentence Encoder) to calculate USE score.
@@
where $N_{\text{words}}$ is the total number of words in the text, $N_{\text{sentences}}$ is the total number of sentences, and $N_{\text{syllables}}$ is the total number of syllables. A smaller $\Delta r$ indicates less change in readability.
@@ A.4.1 Defense Methods
Edit Pretraining (EP) Wang et al. (2024c): An adversarial data augmentation method that blends Mixed Edit Attack into the training set.
Paraphrasing Pretraining (PP) Wang et al. (2024c): An adversarial data augmentation method that blends Paraphrasing Attack into the training set.
CERT-ED Huang et al. (2024): An adversarial data augmentation method that adapts randomized deletion to effectively safeguard natural language classification models from diverse edit-based adversarial operations, including synonym substitution, insertion, and deletion.
RanMask Zeng et al. (2023): An adversarial data augmentation method that randomly masks a proportion of words in the input text, thereby mitigating both word- and character-level adversarial perturbations without assuming prior knowledge of the adversaries’ synonym generation.
Text-RS Zhang et al. (2024b): An adversarial data augmentation method that treats discrete word substitutions as continuous perturbations in the embedding space to reduce the complexity of searching through large discrete vocabularies and bolster the model’s robustness.
Text-CRS Zhang et al. (2024a): An adversarial data augmentation method that is built on randomized smoothing, encompassing various word-level adversarial manipulations—such as synonym substitution, insertion, deletion, and reordering—by modeling them in both embedding and permutation spaces.
@@
RADAR Hu et al. (2023a): An Adversarial Training method that uses a paraphraser (such as DIPPER) as adversary and a fine-tuned model as detector. The adversary and detector learn from each other and improve the robustness of the detector when facing paraphrasing attack.
OUTFOX Koike et al. (2024): An adversarial training method that iteratively improves a detector by generating in-context adversarial examples via a co-evolving attacker. The attacker crafts examples to evade the detector, which are then used to strengthen the detector through in-context learning.
@@ A.5 Detailed Information of Text Perturbation Method
In this section, we introduce 10 Text Perturbation Methods mentioned in Table 1. Detailed information of Adversarial Attack Methods can be found in Appendix A.4.2.
@@
HMGC Zhou et al. (2024): A black-box zero-query attack framework. It uses a surrogate model to approximate the detector, ranks words by gradient sensitivity and PPL, and replaces high-importance words via an encoder-based masked language model. Constraints ensure fluency and semantic consistency, while dynamic adversarial learning refines the attack strategy.
@@
Appendix B Experiment on Defense
@@ B.1 Defense Performance under Different Attack Strengths
B.2 Impact of Synchronous Update of Greater-D
The synchronous update mechanism plays a crucial role in enhancing the robustness of the target detector. To explicitly demonstrate its contribution, we ablate the generator-side updates and instead use static texts to update the GREATER-D detector. The results are presented in Table 7.
Epoch GREATER-D (w/o synchronous update) GREATER-D (w/ synchronous update)
1 59.04 57.64
2 52.17 46.08
3 45.52 21.85
4 40.17 10.28
5 34.52 4.77
6 34.08 3.45
Table 7: Defense performance in ASR (%) of GREATER-D against paraphrasing attacks across different training epochs.
As shown in Table 7, the robustness of the GREATER-D detector significantly degrades when it is trained with static adversarial examples. The ASR converges to approximately 34% after around 5 epochs, which is substantially worse than the result achieved with dynamic updates. This demonstrates the critical role of the dynamic update mechanism in enhancing model robustness.
B.3 Defense Performance of Different Epochs
To demonstrate the effectiveness of the adversarial training procedure, we analyze the GREATER-A and GREATER-D performance after each training round. The detailed analysis of training dynamics is in the Table 8. We present the ASR of GREATER-A and defense performance of GREATER-D against Paraphrasing Attack at each epoch from 1 to 6. We find that as training progresses, the GREATER-A becomes stronger at attacking XLM-Roberta detector. Moreover, GREATER-D exhibits more robust defense performance on the attack it is not trained with.
Epoch GREATER-A ASR GREATER-D ASR
1 77.04 57.64
2 79.03 46.08
3 84.61 21.85
4 88.94 10.28
5 94.06 4.77
6 96.58 3.45
Table 8: ASR (%) of the Adversary against F.t.XLM-RoBERTa-Base detector and the Detector against Paraphrasing Attack across different epochs.
B.4 Experiments on Multilingual data
We evaluate the multilingual performance of different defense methods using German and Urdu datasets Siino (2024), and the results are shown in Table 9. As observed from the table, our model achieves strong performance on both languages and yields the best overall results, indicating its effectiveness in multilingual detection scenarios.
Source Metric F.t.XLM-RoBERTa EP PP VAT TAVAT CERT-ED RanMask Text-RS Text-CRS GREATER-D
German Acc 68.30 59.84 73.06 67.44 59.37 55.60 61.13 65.04 56.80 72.86
F1 75.36 71.13 78.40 75.30 71.08 68.90 71.40 73.51 69.55 77.76
Urdu Acc 56.01 55.09 56.88 63.33 59.04 66.60 64.03 70.62 70.21 66.24
F1 65.27 69.14 69.85 69.07 68.56 74.97 69.81 76.10 76.21 74.89
Overall Acc 62.16 57.47 64.97 65.39 59.21 61.10 62.58 67.83 63.51 69.55
F1 70.32 70.14 74.13 72.19 69.82 71.94 70.61 74.81 72.88 76.33
Table 9: Performance of different defense methods in a multilingual setting. The best result in each group is highlighted with a green background.
Appendix C Evaluations on Greater-A Performance
C.1 Attack on the real-world detector
Speaking of the fact that the adversary and target model share the same training dataset (but different sets), it is necessary to evaluate if the attack conducted by Greater-A would generalize to other detectors which trained on different datasets. To validate this, we attack a close-sourced commercial detector GPTZero [^4] https://gptzero.me/ with Greater-A, which is a strict black-box setting because no other information except for the input and output of the target model is available. The results are in the Table 10. It demonstrates that our Greater-A remains high ASR compared with other SOTA methods, though we are not aware of the architecture, parameters, or training data at all.
Method Avg Queries $\downarrow$ ASR (%) $\uparrow$ Pert. (%) $\downarrow$ $\Delta\mathrm{PPL}$downarrow USE $\uparrow$ $\Delta r$downarrow
HQA 172.38 63 11.23 133.58 0.8312 66.24
FastTextDodger 162.58 70 10.58 96.25 0.9096 42.15
ABP 133.46 82 9.13 72.78 0.8613 36.18
T-PGD 189.36 42 13.12 168.54 0.8077 60.51
Greater-A (ours) 42.57 100 7.82 67.13 0.9393 13.39
Table 10: The attack results of different attack methods on the GPTZero detector. For each sample, the maximum number of queries is limited to 200. The best result in each group is highlighted with a green background.
C.2 Generalization to Other Detector
To demonstrate the generalization of our Greater-A, we conduct additional experiments with two zero-shot detectors: Fast-DetectGPT Bao et al. (2023) and Binoculars Hans et al. (2024). Moreover, to further demonstrate the generalization ability of Greater-A, we also test it on a decoder-only detector F.t.GPT2 Radford et al. (2019) which shares a different architecture with XLM-RoBERTa. As shown in the Table 11, Greater-A outperforms all comparison methods when attacking zero-shot and decoder-only detectors, which demonstrates the universal effectiveness of Greater-A.
Method Fast-DetectGPT Binoculars F.t.GPT2
WordNet 51.36 37.35 42.66
Back Translation 37.50 19.13 1.26
Rewrite 10.97 8.88 30.02
T-PGD 56.83 41.91 70.46
HMGC 17.39 31.43 35.08
Greater-A (ours) 61.14 65.83 74.25
Table 11: The ASR (%) across different detectors under different attack methods. The best result in each group is highlighted with a green background.
Appendix D Ablation Study
@@
Appendix E Mathematical Analysis of Perturbation Rate and Query Complexity
@@
E.1 Perturbation Rate Analysis
@@
E.2 Query Complexity Analysis
@@
Appendix F Adversarial Training Process
@@
Appendix G Case Study
@@
Method Text Result Pert. (%) Queries
Original MGT Suggested by the Scottish Parliamentary Constituencies Commission in 2003, and adopted at Holyrood on 1 May 2004. Area of Scotland’s 32nd largest council area - covering parts of East Renfrewshire Council Area Machine-written [Uncaptioned image] - -
PWWS Suggested by the Scottish Parliamentary Constituencies Commission in 2003, and adoptions at Holyrood on 1 Probability 2004. Area of Scotland’s 32nd highest council area - covering item of East Renfrewshire Council Area Human-written (Succeeded) [Uncaptioned image] 11.76 107
TextFooler Suggested by the Scottish Parliamentary Constituencies Commission in 2003, and adopted at Holyrood on 1 May 2004. Area of Scotland’s 32nd largest council area - covering parts of East Renfrewshire Council Area Machine-written (Failed) [Uncaptioned image] - -
BERTAttack Suggested by the Scottish rural Constituencies Commission in 2003, and abolished at Holyrood on 1 May 2004. Area of ward 32nd most council constituency - almost all of East Renfrewshire Council Area Human-written (Succeeded) [Uncaptioned image] 20.59 60
A2T Suggested by the Scottish Parliamentary Constituencies Commission in 2003, and adopted at Holyrood on 1 May 2004. Area of Scotland’s 32nd largest council area - covering parts of East Renfrewshire Council Area Machine-written (Failed) [Uncaptioned image] - -
FastTextDodger Suggested by the Scottish Parliamentary Constituencies Commission in 2003: and adopted at Holyrood on 1 May 2004. Area of’ ’s 32nd largest council area - covering parts of East council Council Human-written (Succeeded) [Uncaptioned image] 14.71 73
ABP Suggested aside the Scottish Parliamentary Constituencies Commission indium 2003, and adopted atomic number 85 Holyrood on ace Crataegus oxycantha 2004. Area of Scotland’s 32nd largest council area - covering parts of East Renfrewshire Council Area Human-written (Succeeded) [Uncaptioned image] 23.53 71
HQA Suggested by the Scottish Parliamentary Constituencies Commission in 2003, and adopted at Holyrood on 1 May 2004. Area of Scotland’s 2004. sphere council area thirty-second prominent council of East Renfrewshire Council Area Human-written (Succeeded) [Uncaptioned image] 20.59 44
T-PGD Introduced. by. Scottish Parliamentary Resituiances Commission in 2002, and adopted at Holyrood on 1 May 2004. All of Scotland’s 32 The largest councils area - covering parts of East Renfrewshire Council Area Human-written (Succeeded) [Uncaptioned image] 26.47 56
Greater-A Suggested by the Grave Parliamentary Constituencies Commission in 2003, and adopted at Holyrood on 1 May 2004. Area Scotland 32nd biggest council area - covering parts of East Renfrewshire Council Area Human-written (Succeeded) [Uncaptioned image] 5.88 12